# Weakly Supervised Object Segmentation by Background Conditional Divergence

### v1

@@ C.4 Results: MNIST, FasionMNIST, and dSprites on Textures
Figure 8: Background Clusters: From left to right: True clusters, 5, 10, and 15 clusters.
@@
Figure 9: Results of the segmentation model for Fashion MNIST (left), MNIST (middle), and dSprites (right) using true clusters. From top to bottom: input images, estimated masks, true masks, and thresholded estimated masks at .5
@@
C.4.1 Unsupervised Object Segmentation Benchmarks
@@ D.1 SSS Dataset
Figure 12: Segmentation model results on SSS images. Each row corresponds to a different image. From left to right: input images, ground truth masks, estimated masks, and thresholded estimated masks at 0.5. Image sites: Viator.
@@
Figure 13: Segmentation model results on SSS images. Each row corresponds to a different image. From left to right: input images, ground truth masks, estimated masks, and thresholded estimated masks at 0.5. Images sites from top to bottom: Barge, WH Gilbert , WH Gilbert, and Barge.
Figure 14 shows images without ships, where the model identifies the artificial reef, although it not in the ground truth. As a man-made object and arrangement it could be argued that the masking network worked in this case. However, it also highlights an image where there are extensive false positives for portions of the seafloor.
Refer to captionRefer to caption
Figure 14: Segmentation model results on SSS images without ships, but regular arrangement of rocks or blocks in the first image. Each row corresponds to a different image. From left to right: input images, ground truth masks, estimated masks, and thresholded estimated masks at 0.5. Site of the first image is Artificial Reef, and for the second image is Haltiner Berge.
@@
Figure 15: Failure cases of the segmentation model on SSS images at ‘James Davidson’. From left to right: input image, ground truth mask, estimated mask, and thresholded estimated mask at 0.5.
D.2 Benchmark Results
Attempts to apply unsupervised segmentation methods to the SSS dataset are shown in Figures 16, 17, 18, and 19.
Figure 16: Two blocks of results from Chen et al. (2019) model on the SSS dataset. Each row represents the result of an input sonar patch: first column is the input sonar patch, second is the segmentation ground truth, third and fourth are generated masks, and the last two are alpha blended results based on the generated masks.
Figure 17: Two blocks of results from Chen et al. (2019) model on the SSS dataset. Each row represents the result of an input sonar patch: first column is the input sonar patch, second is the segmentation ground truth, third and fourth are generated masks, and the last two are alpha blended results based on the generated masks. The left figure shows a group of background patches and the right one shows a group of composite images.
Figure 18: Generated images from ILSGAN (Zou et al., 2023) on the SSS dataset. Each cell represents a generated image.
Figure 19: Results from model-free unsupervised baseline method  Savarese et al. (2021) using patches with 50% overlaps on the SSS dataset.
D.3 Additional Results for CUB Dataset
Additional counterfactuals from the model trained on the original CUB split are shown in Figure 20.
Figure 21 shows examples of the background images extracted from original CUB split organized into 10 clusters. Note that color and color are very distinct.
@@ D.3 Additional Results for CUB Dataset
Sensitivity study of performance metrics for the choice of latent representation (auto-encoder or contrastive learning) and the number of clusters for the original CUB split are shown in Table 6.
Table 6: Masking Network Performance (Average Precision, AUCROC and IoC) Comparison of AE+K-means and CL+K-Means based Background Clustering

### v2

@@ C.3 Additional Results for Textures+MNIST/FashionMNIST/dSprites
Figure 10: Background Clusters for Textures: From left to right: True clusters, 5, 10, and 15 clusters.
Figure 11 shows qualitative results for the masks using the true texture classes.
@@
Figure 11: Results of the segmentation model for Texture+Fashion MNIST (left), Texture+MNIST (middle), and Texture+dSprites (right) using true clusters. From top to bottom: input images, estimated masks, true masks, and thresholded estimated masks at 0.5
@@
C.4 Unsupervised Object Segmentation Benchmarks for Textures+MNIST/FashionMNIST/dSprites
@@ D.1 Additional Results for AI4Shipwrecks
Figure 14: Segmentation model results on AI4Shipwrecks. Each row corresponds to a different image. From left to right: input images, ground truth masks, estimated masks, and thresholded estimated masks at 0.5. Image sites: Viator.
@@
Refer to captionRefer to caption
Figure 15: Segmentation model results on images without ships from AI4Shipwrecks, but regular arrangement of rocks or blocks in the first image. Each row corresponds to a different image. From left to right: input images, ground truth masks, estimated masks, and thresholded estimated masks at 0.5. Site of the first image is Artificial Reef, and for the second image is Haltiner Berge.
@@
Figure 16: Segmentation model results on AI4Shipwrecks. Each row corresponds to a different image. From left to right: input images, ground truth masks, estimated masks, and thresholded estimated masks at 0.5. Images sites from top to bottom: Barge, WH Gilbert , WH Gilbert, and Barge.
Figure 15 shows images without ships, where the model identifies the artificial reef, although it not in the ground truth. As a man-made object and arrangement it could be argued that the masking network worked in this case. However, Figure 15 also highlights an image where there are extensive false positives for portions of the seafloor.
@@
Figure 17: Failure cases of the segmentation model on images at ‘James Davidson’ from AI4Shipwrecks. From left to right: input image, ground truth mask, estimated mask, and thresholded estimated mask at 0.5.
D.1.1 Unsupervised Benchmark Results on AI4Shipwrecks
Attempts to apply unsupervised segmentation methods to the AI4Shipwrecks dataset are shown in Figures 18, 19, 20, and 21.
Figure 18: Two blocks of results from Chen et al. (2019) model for AI4Shipwrecks. Each row represents the result of an input sonar patch: first column is the input sonar patch, second is the segmentation ground truth, third and fourth are generated masks, and the last two are alpha blended results based on the generated masks.
Figure 19: Two blocks of results from Chen et al. (2019) model for AI4Shipwrecks. Each row represents the result of an input sonar patch: first column is the input sonar patch, second is the segmentation ground truth, third and fourth are generated masks, and the last two are alpha blended results based on the generated masks. The left figure shows a group of background patches and the right one shows a group of composite images.
Figure 20: Generated images from ILSGAN (Zou et al., 2023) for AI4Shipwrecks. Each cell represents a generated image.
Figure 21: Results from model-free unsupervised baseline method Savarese et al. (2021) using patches with 50% overlaps on AI4Shipwrecks.
D.2 Additional Results for CUB Dataset
Figure 22 shows examples of the background images extracted from original CUB split organized into 10 clusters. Note that color and texture are very distinct.
@@ D.2 Additional Results for CUB Dataset
For quantitative results, we show a sensitivity analysis of performance metrics for the choice of latent representation (auto-encoder or contrastive learning) and the number of clusters for the original CUB split are shown in Table 7.
Table 7: Masking Network Performance (Average Precision, AUCROC and IoU) Comparison of AE+K-means and CL+K-Means based Background Clustering